A Complete Bibliography of Scientific American: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.08

Title word cross-reference

100-Billion-Volt [?]. 1900-1950 [?].

Absorption [?]. Abundance [?]. Abyss [?, ?]. Accelerator [?, ?]. Accelerators [?]. Acids [?, ?]. Acquired [?]. Act [?, ?]. ACTH [?]. Action [?]. Activity [?]. Addendum [?]. Aerodynamics [?, ?]. Aether [?]. Africa [?]. against [?]. Agamemnon [?]. Agammaglobulinemia [?]. Age [?, ?, ?, ?]. Aging [?]. Ago [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Ago [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Agora [?]. Agricolae [?]. Agriculture [?, ?]. Air [?]. Airglow [?]. Airplanes [?]. Airport [?]. Albert [?]. Alchemists [?]. Alchemy [?]. Alcohol [?]. Aleuts [?]. Alfred [?]. Algae [?, ?]. Aligned [?]. Alkaloids [?]. Almanac [?]. Alone [?]. Altitude [?]. Altitudes [?]. Amateur [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Amateur [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Amber [?]. Ambo [?]. Ambroise [?]. America [?, ?]. American [?, ?]. Americans [?]. Amoebae [?]. Amplifiers [?]. Analysis [?, ?, ?]. Anatomy [?]. Ancient [?, ?, ?, ?, ?]. Anemia [?]. Anesthesia [?]. Anger [?]. Angiotensin [?]. Animal [?, ?, ?, ?, ?, ?]. Animals [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Annual [?, ?, ?, ?, ?, ?, ?]. Ant [?]. Anthropology [?, ?, ?, ?]. Anti [?, ?]. Anti-Matter [?]. Anti-Scientific [?]. Antibiotics [?, ?]. Antibodies [?, ?, ?]. Antiochus [?]. Antiproton [?]. Antiquity [?, ?]. Appetite [?]. Applied [?]. appreciation [?]. Archaeology [?, ?]. Arctic [?, ?, ?]. Aristotle [?]. Arms [?]. Armstrong [?]. Arrested [?]. Arroyos [?]. Art [?, ?, ?]. Articles [?, ?, ?]. Artificial [?, ?, ?, ?, ?, ?, ?]. Asthma [?]. Astronomer [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Astronomy [?, ?, ?]. Atka [?]. Atlantic [?]. Atmosphere [?, ?, ?, ?, ?, ?, ?]. Atom [?, ?, ?]. Atomic [?, ?, ?, ?, ?, ?, ?, ?, ?]. Atoms [?, ?]. Atp [?]. Attitude [?]. Attitudes [?, ?]. Audubon [?]. Aurora [?]. Authors [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Autobiographies [?]. Automatic [?, ?, ?, ?]. Autumn [?]. Avalanches [?]. Aviation [?].

Babbage [?]. Babies [?]. Bacteria [?, ?, ?, ?, ?]. Bacterial [?]. Balance [?]. Ballistocardiography [?]. Balloon [?]. Barbiturates [?]. Barrier [?]. Barriers [?]. Basis [?]. Bat [?]. Bathyscaph [?]. Bats [?, ?]. Battery [?]. Beer [?]. Bees [?]. Behavior [?, ?, ?]. Belts [?]. Berylliosis [?]. Beryllium [?]. Bevatron [?]. Beyond [?]. Bibliography [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Bibliography [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Big [?]. Billion [?]. Biochemistry [?]. Biological [?, ?]. Biology [?, ?, ?]. Bird [?, ?]. Birds [?, ?, ?, ?, ?, ?, ?]. Birth [?]. Blackjack [?]. Blast [?]. Bligh [?]. Blight [?]. Blind [?]. Blindness [?]. Blood [?, ?, ?, ?, ?, ?, ?, ?]. Blue [?]. Boards [?]. Body [?, ?, ?, ?, ?, ?]. Bog [?]. Bogs [?]. Boiling [?]. Bomb [?, ?, ?, ?, ?]. Bone [?]. Books [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Books [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Borders [?]. Boredom [?]. Bottom [?]. Boundary [?]. Bourbaki [?]. Bower [?]. Boy [?]. Brain [?, ?, ?, ?, ?]. Brains [?]. Breadfruit [?]. Breathing [?]. Breeder [?]. Bridges [?, ?]. Britain [?]. Bubble [?]. Bubbles [?]. Builders [?]. Buildings [?]. Burial [?]. Business [?].

C [?]. Calcium [?]. Calls [?]. came [?]. Camel [?]. Camera [?]. Camp [?]. Canadian [?]. Canary [?]. Cancer [?, ?, ?]. Canyon [?]. Canyons [?]. Captain [?]. Captivity [?]. Carbon [?]. Cardiogram [?]. Cargo [?]. Carlo [?]. Carotenoids [?]. Carroll [?]. Case [?, ?]. Castles [?, ?]. Cataclysmic [?]. Catastrophe [?]. Cats [?]. Cattle [?]. Cave [?, ?]. Caves [?]. Celestial [?]. Cell [?, ?, ?, ?, ?, ?, ?]. Cells [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Cellulose [?]. Census [?]. Centered [?]. Centers [?]. century [?]. Cerebellum [?]. Challenger [?]. Chamber [?]. Change [?]. Changing [?, ?, ?]. Characteristics [?]. Charles [?, ?, ?, ?]. Chelation [?]. Chemical [?, ?, ?, ?, ?, ?, ?]. Chemistry [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Chess [?, ?]. Chess-Player [?]. Chess-Playing [?]. Chick [?]. Chickens [?]. Child [?]. Children [?, ?, ?, ?, ?, ?, ?]. Chromatography [?]. Chromosomes [?]. Circuits [?]. Circulation [?, ?, ?, ?, ?]. Circulatory [?]. Cities [?, ?]. Citizen [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Citizen [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. City [?, ?, ?]. Civilization [?]. Clearance [?]. Clerk [?]. Client [?]. Client-Centered [?]. Clifford [?]. Climate [?, ?, ?, ?, ?]. Clocks [?, ?]. Clothes [?]. Clothing [?]. Clots [?]. Cloud [?]. Clouds [?, ?, ?]. Coal [?, ?]. Cocoons [?]. Coelacanth [?]. Cold [?, ?]. Colliding [?]. Colony [?]. Color [?, ?, ?, ?, ?]. Colors [?]. Combat [?]. come [?]. Comets [?, ?]. Common [?]. Community [?]. Complex [?]. Compression [?]. Computer [?, ?, ?, ?, ?]. Computers [?, ?]. Concepts [?]. Concrete [?]. Conditioners [?]. Conditioning [?, ?]. Conference [?]. Conferences [?]. Conflict [?]. Congenital [?]. Content [?]. Continental [?]. Continents [?]. Contraction [?]. Control [?, ?, ?, ?, ?, ?, ?, ?, ?]. Cooperation [?]. Cord [?]. Coriolis [?]. Corn [?, ?]. Coronary [?]. Corpuscles [?]. Corrosion [?]. Cortisone [?]. Cosmic [?, ?, ?]. Cosmology [?, ?]. Counters [?, ?]. Courtship [?, ?]. Crab [?, ?]. Crater [?]. Craters [?]. Creative [?]. Crisis [?]. Crop [?]. Crows [?]. Crust [?]. Crystals [?, ?, ?]. Cults [?]. Cultural [?]. Culture [?, ?]. Cultures [?, ?]. Curiosity [?]. Curious [?]. Current [?]. Curtain [?]. Curvature [?]. Cutting [?]. Cycle [?].

D’Arcy [?]. Darwin [?, ?, ?]. Darwinism [?]. Dating [?]. Death [?, ?, ?, ?]. Decay [?]. Decisions [?]. Deep [?]. Deep-Sea [?]. Deer [?]. Defects [?]. Defense [?]. Deformities [?]. Depths [?]. Descartes [?]. Desegregation [?]. Desert [?, ?, ?]. Deserts [?]. Destruction [?]. Detergents [?]. Development [?]. Diamonds [?]. Different [?]. Differentiate [?]. Differentiation [?, ?, ?]. Diffusion [?]. Dig [?]. Diode [?]. Dioxide [?]. Diphtheria [?]. Disclosure [?]. Discoveries [?]. Discovery [?]. Discrimination [?]. Disease [?, ?, ?, ?]. Diseases [?]. Dislocations [?]. Dissociated [?]. Distances [?]. Distribution [?]. Division [?]. do [?, ?, ?, ?]. Doctors [?]. Dog [?]. Dogs [?]. Dream [?]. Drilling [?]. Drugs [?, ?, ?, ?]. Dune [?]. Dunkers [?]. Duplication [?]. Dust [?]. Dying [?, ?]. Dynamo [?].

Ear [?]. Earliest [?]. Early [?, ?, ?, ?]. Ears [?]. Earth [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Earthquake [?]. Earths [?]. Easter [?]. Eclipses [?]. Ecology [?]. Economic [?]. Economics [?, ?]. Ecosphere [?]. Edible [?]. Edmund [?]. Edwin [?]. Eelgrass [?]. Effect [?]. Effects [?, ?]. Egg [?]. Einstein [?, ?]. Election [?]. Electoral [?]. Electric [?]. Electrical [?, ?]. Electricity [?, ?, ?, ?]. Electroluminescence [?]. Electronic [?, ?]. Electronics [?, ?]. Electrophoresis [?]. Element [?]. Elementary [?, ?, ?]. Elements [?, ?, ?, ?, ?, ?, ?]. Elephants [?]. Embryo [?]. Embryologist [?]. Eminent [?]. Emotions [?]. Empire [?]. Empty [?]. Encouragement [?]. End [?]. Energy [?, ?, ?]. Engineers [?]. Engines [?]. Enteroviruses [?]. Environment [?]. Enzyme [?]. Enzyme-Substrate [?]. Epidemiology [?]. Equipment [?]. Era [?]. Eradication [?]. Ernest [?]. Errata [?, ?, ?, ?, ?, ?, ?, ?]. Erratum [?, ?, ?, ?, ?, ?]. Essential [?]. Estuary [?]. Etruscan [?]. Euler [?]. European [?]. Events [?]. Evidence [?]. Evolution [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Evolutionary [?]. Examines [?]. Exchange [?]. Exclusion [?]. Executive [?]. Experiment [?]. Experimental [?, ?]. Experiments [?, ?, ?, ?, ?, ?, ?, ?]. Explanation [?]. Exploring [?]. Explosives [?]. Eye [?, ?, ?].

F [?]. Fall [?]. Family [?, ?]. Faraday [?]. Farmer [?]. Fat [?]. Father [?]. Fats [?]. Fear [?]. Feedback [?, ?]. Fertility [?, ?]. Fertilization [?, ?, ?, ?, ?]. Fever [?]. Fibers [?]. Fiddler [?]. Field [?]. Fields [?]. Filariasis [?]. Finches [?]. Fine [?]. Fire [?, ?]. Fire-Maker [?]. First [?, ?]. Fishes [?]. Fission [?, ?]. FitzGerald [?]. Fixation [?]. Flagella [?]. Flame [?]. Flares [?]. Flexagons [?]. Flight [?, ?, ?, ?]. Floor [?, ?, ?]. Flow [?]. Flowering [?, ?]. Flowers [?, ?]. Flying [?, ?]. Flying-Spot [?]. Food [?, ?]. Force [?]. Forecasting [?]. Forest [?]. Forgotten [?, ?, ?]. Form [?, ?, ?]. Formation [?]. Fort [?]. Fossil [?, ?]. Foundation [?, ?]. Four [?]. Fractures [?]. Francis [?, ?]. Free [?, ?]. Freezes [?]. Freezing [?]. Fresh [?]. Friction [?]. Frog [?]. Frostbite [?]. Frozen [?]. Fruit [?].

Fuel [?, ?, ?]. Fundamental [?]. Fungi [?]. Fusion [?]. Future [?, ?].

G. [?]. Galaxies [?, ?, ?, ?, ?, ?]. Galaxy [?, ?, ?]. Galen [?]. Galls [?]. Galton [?]. Game [?]. Games [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Gamma [?]. Gas [?, ?, ?, ?]. Gene [?]. General [?]. Generalized [?]. Genes [?, ?]. Genetic [?, ?, ?]. Genetics [?, ?, ?, ?]. Geneva [?, ?, ?, ?]. Genius [?]. Geography [?, ?, ?]. Geological [?]. Geology [?]. Geometry [?, ?]. Georgii [?]. Germ [?, ?]. Germination [?]. Giant [?, ?, ?, ?]. Glaciers [?, ?]. Glands [?]. Glaucoma [?]. Globulin [?]. Gödel [?]. Gomera [?]. Gorge [?]. Gout [?]. Grand [?]. Granite [?]. Grasshopper [?]. Gravitation [?]. Gravity [?]. Gray [?]. Great [?]. Greek [?]. Ground [?]. Group [?, ?]. Groups [?, ?]. Growth [?, ?, ?, ?, ?].

H [?]. half [?]. half-century [?]. Halloween [?]. Hamilton [?]. Harvey [?]. Hauksbee [?]. Health [?]. Heart [?, ?, ?, ?, ?]. Heartbeats [?]. Heat [?, ?, ?, ?, ?, ?, ?]. Heavy [?]. Heinrich [?]. Helicopters [?]. Helmholtz [?]. Henry [?]. here [?]. Hereditary [?, ?]. Heredity [?, ?]. Heroes [?]. Hertz [?]. Hibernation [?]. High [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. High-Energy [?]. High-Speed [?, ?]. Highlands [?]. History [?, ?, ?, ?, ?]. Holds [?]. Home [?]. Homer [?]. Homing [?]. Homunculus [?]. Honeybee [?]. Hoo [?]. Hooke [?]. Hopi [?]. Hormone [?]. Hormones [?]. Horn [?]. Hot [?, ?]. House [?]. Human [?, ?, ?, ?, ?, ?, ?, ?]. Hummingbirds [?]. Humor [?]. Hunter [?]. Hurricanes [?, ?]. Hutterites [?]. Hybrid [?]. Hydra [?, ?]. Hydrazine [?]. Hydrogen [?, ?, ?, ?, ?]. Hypnosis [?]. Hypothermia [?].

Ice [?]. Idea [?]. Ideas [?]. II [?, ?]. III [?]. Imagination [?, ?]. Imitation [?]. Imitative [?]. Imprinting [?]. Impulse [?, ?]. Incriminating [?]. Incubator [?]. Indestructible [?]. Index [?, ?, ?, ?, ?, ?, ?, ?, ?]. Indo [?]. Indo-European [?]. Industrial [?]. Inertia [?]. Infant [?, ?]. Infinity [?]. Influence [?]. Influenza [?, ?]. Information [?, ?, ?, ?]. Infrared [?]. Inherited [?]. Innovation [?, ?, ?, ?]. Input [?]. Input-Output [?]. Insect [?, ?, ?, ?, ?]. Insecticides [?, ?]. Insects [?, ?, ?, ?]. Instincts [?]. Instrument [?, ?, ?]. Instruments [?]. Insulin [?, ?]. Intellectual [?]. Intelligence [?, ?]. Interact [?]. Interactions [?]. Interior [?]. Internal [?]. International [?]. Interstellar [?]. Interview [?]. Intuition [?]. Invention [?]. Ion [?]. Ionizing [?, ?, ?, ?, ?, ?, ?, ?]. Ionosphere [?]. Isaac [?]. Island [?]. Islands [?]. Isotopes [?]. IV [?].

James [?]. Jericho [?]. Jet [?, ?]. Jewish [?]. Jocular [?]. Joey [?]. Joseph [?]. Junction [?, ?]. Junction-Diode [?]. Jupiter [?]. Juvenile [?].

Kidney [?, ?]. King [?]. Kingdon [?]. Kinship [?]. Klystron [?].

Koenigsberg [?]. Kuanyama [?]. Kwashiorkor [?].

Labor [?, ?]. Laboratories [?]. Lamont [?]. Lamprey [?]. Language [?, ?, ?, ?, ?, ?, ?, ?]. Languages [?, ?]. Laplace [?]. Larger [?]. Late [?, ?]. Lauritsen [?]. Lavoisier [?]. Lawn [?]. Laws [?]. Layer [?, ?]. Leap [?]. Learning [?, ?]. Learns [?]. Leaves [?]. Leonhard [?]. Lethal [?, ?]. Letters [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Letters [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Lewis [?]. Lichens [?]. Life [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Light [?, ?, ?, ?, ?, ?]. Lignin [?]. Limits [?]. Line [?]. Linear [?, ?]. Liquids [?]. Living [?, ?, ?, ?, ?, ?]. Lobotomy [?]. Locusts [?]. Logic [?, ?]. Long [?, ?]. Long-Range [?]. Lost [?]. Love [?]. Low [?]. Low-Speed [?]. Lyell [?].

Machine [?, ?, ?, ?, ?, ?]. Machines [?, ?, ?, ?, ?]. Made [?, ?, ?]. Magellan [?]. Magnetic [?, ?, ?]. Magnetism [?, ?]. Maker [?]. Makes [?]. Making [?]. Malaria [?]. Male [?]. Mammals [?]. Man [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Manners [?]. Manufacture [?]. Mars [?]. Maser [?]. Mass [?, ?]. Masters [?]. Material [?]. Materials [?, ?]. Mathematical [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mathematician [?]. Mathematics [?, ?, ?]. Matter [?, ?, ?, ?]. Maupertuis [?]. Maxwell [?, ?]. May [?]. Maya [?]. Measured [?]. Measurement [?]. Measuring [?]. Mechanical [?, ?]. Medicine [?, ?]. Melody [?]. Memories [?]. Memory [?, ?]. Men [?, ?, ?]. Mental [?, ?]. Mesonic [?]. Messengers [?]. Metabolism [?, ?, ?, ?, ?]. Metallica [?]. Metallurgy [?]. Metals [?, ?, ?, ?, ?]. Metamorphosis [?, ?]. Meteor [?, ?]. Meteorite [?]. Meteorites [?]. Meteors [?]. Method [?]. Methods [?, ?]. Metropolitan [?]. Mice [?]. Michael [?]. Microcirculation [?]. Micrographs [?]. Microscope [?, ?, ?, ?]. Microsome [?]. Microsurgery [?]. Microwaves [?]. Midas [?]. Military [?]. Milk [?]. Milky [?]. Milling [?]. Mind [?]. Mine [?]. Missing [?]. Mississippi [?, ?]. Moas [?]. Mobilization [?]. Model [?]. Modern [?, ?]. Mohole [?]. Molds [?]. Molecular [?, ?]. Molecule [?, ?]. Molecules [?, ?, ?, ?, ?]. Moment [?]. Mongolism [?]. Monkey [?]. Monkeys [?, ?, ?]. Monmouth [?]. Monsters [?]. Monte [?]. Moon [?]. Mortality [?, ?]. Motion [?]. Motions [?]. Mound [?]. Movements [?]. Moving [?]. Multiplication [?]. Multiplicity [?]. Muscle [?, ?, ?, ?]. Museum [?]. Mushrooms [?]. Music [?]. Musical [?]. Mutation [?, ?]. Mycenae [?]. Mystery [?, ?].

Nation [?]. National [?, ?, ?]. Natural [?, ?]. Nature [?, ?, ?, ?]. Navigation [?, ?, ?, ?]. Neanderthal [?]. Nebula [?]. Negro [?]. Neptune [?]. Nerve [?, ?, ?]. Nervous [?]. Nestor [?]. Net [?]. Neuroses [?]. Neutrino [?]. Neutron [?]. Newest [?]. Newton [?]. Nicolas [?]. Nitrogen [?]. Nocturnal [?]. Non [?]. Non-Military [?]. Norman [?]. North [?]. Norway [?]. Nuclear [?, ?, ?, ?, ?, ?]. Nuclei [?]. Nucleic [?, ?]. Nucleus [?, ?, ?, ?, ?]. Numbers [?, ?].

Oaks [?]. Obelisk [?]. Obesity [?]. Observations [?]. Observatory [?, ?]. Observed [?]. Obstetrical [?]. Ocean [?, ?]. Oceans [?]. Oil [?]. Oils [?]. Older [?]. Oldest [?, ?]. Olduvai [?]. Operations [?]. Opinions [?, ?]. Opossum [?]. Order [?]. Oreopithecus [?]. Organic [?, ?]. Organizer [?]. Organs [?]. Origin [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Origins [?]. Output [?]. Overthrow [?]. Oxidation [?]. Oysters [?].

Pacific [?, ?, ?]. Pain [?]. Palace [?]. Paleobiochemistry [?]. Palomar [?, ?]. Panel [?]. Papyrus [?]. Paradoxes [?]. Paralytic [?]. Paré [?]. Parentage [?]. Parity [?]. Particle [?]. Particles [?, ?, ?, ?, ?]. Partner [?]. Parts [?]. Pathology [?]. Peat [?]. Penguins [?]. People [?, ?, ?]. Peoples [?]. Perception [?, ?, ?, ?, ?]. Perfect [?]. Perkin [?]. Personality [?]. Peru [?, ?]. Peruvian [?]. Petroleum [?]. pH [?]. Pharaohs [?]. Phosphors [?]. Photocell [?]. Photographic [?]. Photography [?]. Photosynthesis [?]. Physicist [?]. Physics [?, ?, ?, ?, ?, ?, ?, ?, ?]. Physiology [?, ?, ?, ?]. Pidgin [?]. Pigments [?]. Pile [?]. Pine [?]. Pions [?]. Pituitary [?]. Placebos [?]. Plague [?, ?]. Planets [?]. Plankton [?]. Planning [?]. Plant [?, ?, ?, ?, ?, ?, ?, ?]. Plants [?, ?, ?, ?]. Plasma [?]. Plasmoids [?]. Player [?]. Playing [?, ?]. Pleasure [?]. Pneumatic [?]. Poetry [?]. Point [?]. Poisonous [?]. Poisons [?, ?]. Poker [?]. Polarized [?]. Policy [?]. Polio [?, ?]. Poliomyelitis [?]. Pollution [?, ?]. Polyethylene [?]. Polymers [?]. Polymorphism [?]. Polynesia [?]. Pompeii [?]. Pond [?]. Population [?, ?, ?]. Populations [?, ?]. ’Possum [?]. Posture [?]. Potato [?]. Potatoes [?]. Power [?, ?, ?, ?, ?, ?, ?]. Powerhouse [?]. Practice [?]. Prairie [?]. Predatory [?]. Prefrontal [?]. Prehistoric [?]. Prejudice [?]. Pressure [?, ?, ?]. Pressures [?]. Prestressed [?]. Priestley [?]. Primitive [?]. Principle [?, ?]. Probability [?, ?, ?, ?]. Problem [?, ?]. Process [?, ?]. Productivity [?]. Products [?]. Progesterone [?]. Programming [?]. Programs [?]. Progress [?, ?, ?]. Project [?]. Projective [?]. Proof [?]. Properties [?]. Propulsion [?]. Prospecting [?]. Protein [?, ?, ?]. Proteins [?, ?, ?]. Proto [?]. Proto-Castles [?]. Protozoon [?]. Psychiatric [?]. Psychologist [?]. Psychology [?, ?, ?, ?]. Psychoses [?]. Psychotherapy [?, ?, ?]. Pulsating [?]. Pump [?]. Pure [?]. Purple [?]. Pursuit [?].

Quality [?]. Quantum [?]. Questions [?]. Quicksand [?].

Rabbit [?]. Races [?]. Radar [?, ?, ?]. Radiation [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Radicals [?, ?]. Radio [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Radioactive [?, ?, ?]. Radiocarbon [?]. Rain [?, ?]. Raindrops [?]. Range [?]. Rare [?]. Rat [?]. Rattlesnake [?]. Rays [?, ?]. re [?]. Reactions [?]. Reactor [?, ?, ?, ?, ?]. Reactors [?, ?, ?]. Really [?]. reasonable [?]. Rebuilding [?]. Records [?]. Red [?, ?, ?]. Red-Shift [?]. Regeneration [?]. Regulate [?]. Relationships [?]. Relativity [?]. Repetition [?]. Reproducing [?, ?]. Reptiles [?]. Research [?, ?, ?, ?, ?]. Resonance [?]. Resources [?, ?, ?]. Respiration [?]. Responses [?]. Reticular [?]. Return [?]. Reversal [?]. Review [?, ?]. Revival [?, ?]. Revolution [?]. Rh [?]. Rhind [?]. Rickettsiae [?]. Rings [?]. Ripening [?]. Rise [?]. River [?]. Roads [?]. Robert [?]. Rocket [?, ?]. Rockets [?]. Role [?]. Rome [?]. Rowan [?]. Rubber [?]. Ruling [?]. Ruminants [?]. Running [?]. Russel [?].

Salmon [?]. Salt [?, ?, ?]. Sand [?]. Sardinia [?]. Sargasso [?]. Satellite [?, ?, ?]. Satellites [?, ?, ?]. Scars [?]. Scattered [?]. Schizophrenia [?, ?]. Schizophrenic [?]. Schmidt [?]. Science [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Science [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Scientific [?]. Scientist [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Scientists [?, ?, ?, ?]. Scintillation [?]. Scopes [?]. Sea [?, ?, ?, ?]. Second [?]. Seeding [?]. Segregation [?]. Selection [?]. Self [?, ?, ?]. Self-Disclosure [?]. Self-Reproducing [?, ?]. Sense [?]. Senses [?]. Separating [?]. Serotonin [?]. Settlement [?]. Sex [?, ?]. Sexuality [?]. Shakespeare [?]. Shale [?]. Shanidar [?]. Shape [?, ?]. Sharks [?]. Shelf [?]. Sherrington [?]. Shift [?]. Ship [?]. Ship-Burial [?]. Shock [?, ?]. Shortest [?]. Shrews [?]. Sickle [?, ?]. Sickle-Cell [?]. Sieves [?, ?]. Simon [?]. Simple [?]. Single [?]. Sir [?, ?]. Situation [?]. Skin [?, ?]. Sky [?, ?]. Sleep [?]. Slip [?]. Slow [?]. Smell [?]. Smog [?]. Snail [?]. Snakebite [?]. Snow [?]. Soap [?]. Social [?, ?, ?, ?, ?, ?]. Soil [?, ?]. Solar [?, ?, ?, ?, ?, ?]. Solids [?, ?, ?]. Sonar [?]. Sounds [?, ?]. Southern [?]. Space [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Specificity [?]. Spectrometer [?]. Speed [?, ?, ?, ?]. Sperm [?]. Spider [?, ?]. Spinal [?]. Spiral [?, ?]. Spot [?]. Spots [?]. Squid [?, ?]. Stains [?]. Staphylococcus [?]. Starlight [?]. Starling [?]. Stars [?, ?, ?, ?, ?, ?, ?]. State [?, ?]. Statistics [?]. Stay [?]. Steady [?]. Steady-State [?]. Steel [?]. Stellar [?]. Stellarator [?].

Steroids [?]. Stick [?]. Stickleback [?]. Stitching [?]. Stock [?]. Stone [?, ?]. Stonehenge [?]. Storms [?]. Straight [?]. Strange [?, ?]. Strangler [?]. Stream [?, ?]. Stress [?]. Strikes [?]. Strong [?]. Structure [?, ?, ?, ?, ?, ?]. Studies [?]. Study [?, ?, ?, ?, ?]. Subjective [?]. Submarine [?]. Substances [?]. Substrate [?]. Suicide [?]. Sumerian [?]. Sumerians [?]. Sun [?, ?, ?, ?, ?, ?, ?, ?]. Sunspots [?]. Superconductivity [?]. Superfluidity [?]. Supergalaxy [?]. Surgery [?]. Surgical [?, ?]. Sutton [?]. Swift [?]. Swim [?]. Switch [?]. Symbolic [?]. Synthesis [?, ?]. Synthetic [?, ?, ?, ?, ?]. System [?, ?, ?, ?]. Systems [?].

T2 [?]. Tails [?]. Takes [?]. Talking [?]. Taste [?]. Teach [?]. Teaching [?, ?, ?]. Technology [?]. Teenage [?]. Teeth [?]. Telescopes [?]. Television [?, ?, ?, ?]. Temperature [?]. Temperatures [?, ?, ?, ?, ?, ?, ?]. Termite [?]. Tewa [?]. their [?]. Theories [?]. Theory [?, ?, ?, ?, ?, ?, ?, ?]. There [?]. Thermoelectricity [?]. Things [?]. Thirst [?]. Thompson [?]. Thoughts [?]. Thrombosis [?]. Thunderstorm [?]. Tides [?, ?]. Time [?]. Tissue [?, ?]. Tissues [?, ?]. Together [?]. Tomb [?]. Tombs [?]. Tones [?]. Too [?]. Tool [?]. Tooth [?]. Topographic [?]. Topology [?]. Tornadoes [?]. Toxin [?]. Toxoplasmosis [?]. Trace [?, ?]. Trace-Element [?]. Tracking [?]. Tracks [?]. Trade [?, ?]. Trade-Wind [?]. Transduction [?]. Transfer [?, ?]. Transformed [?]. Transistor [?]. Translation [?]. Transmission [?]. Transplantation [?]. Transplanting [?]. Transplants [?]. Traumatic [?]. Tree [?]. Trees [?]. Trenches [?]. Trial [?]. Tritium [?]. Trout [?]. Tsunamis [?]. Tuberculosis [?, ?]. Turbine [?]. Turbulence [?]. Turkey [?]. Twins [?].

U.S. [?, ?, ?, ?]. U.S.S.R. [?]. Ulcers [?]. Ultimate [?]. Ultracentrifuge [?]. Ultrahigh [?, ?, ?]. Ultrahigh-Altitude [?]. Ultramicrochemistry [?]. Ultrasonics [?]. Umbilical [?]. Uncertainty [?]. Underwater [?, ?]. Unicorn [?]. Universal [?]. Universe [?, ?, ?, ?, ?]. Unknown [?]. Unorthodox [?]. unreasonable [?]. Upright [?]. Uranium [?, ?]. Useful [?]. Uses [?, ?].

v [?, ?, ?, ?, ?]. Vaccines [?]. Vacuum [?]. Valley [?, ?]. Values [?]. Vanishing [?]. Versatile [?]. Very [?]. Viewed [?]. Village [?]. Virus [?, ?, ?, ?, ?]. Viruses [?, ?, ?, ?, ?, ?]. Vision [?, ?, ?, ?]. Visual [?]. Visualized [?]. Vitro [?]. Volcanoes [?, ?]. Volt [?]. Volvox [?]. Votes [?]. Voyage [?, ?].

Walk [?]. Wall [?]. Wallace [?]. Warm [?]. Warning [?]. Wasp [?]. Water [?, ?, ?, ?, ?, ?]. Waves [?, ?, ?, ?, ?, ?, ?, ?]. Way [?]. Weak [?]. Weather [?, ?, ?]. Webs [?]. Weed [?]. Whale [?, ?, ?]. Whales [?]. Wheat [?]. Where [?]. Whistled [?]. Whistlers [?]. White [?]. Whittaker [?]. Whole [?]. Wild [?]. William [?, ?, ?, ?]. Wind [?].

Windows [?]. within [?]. Witness [?]. Women [?]. Wonderful [?]. Wood [?]. Woodroach [?]. World [?, ?, ?, ?]. Worm [?]. Wound [?].

Years [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Years [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Yerkes [?]. Young [?]. Youth [?].

Zirconium [?].